# Towards Locally Deployable Fine-Tuned Causal Large Language Models for Mode Choice Behaviour

### v1

@@ Abstract
This study investigates the adoption of open-access, locally deployable causal large language models (LLMs) for travel mode choice prediction and introduces LiTransMC, the first fine-tuned causal LLM developed for this task. We systematically benchmark eleven LLMs (1-12B parameters) across three stated and revealed preference datasets, testing 396 configurations and generating over 79,000 synthetic commuter predictions. Beyond predictive accuracy, we evaluate models generated reasoning using BERTopic for topic modelling and a novel Explanation Strength Index, providing the first structured analysis of how LLMs articulate decision factors in alignment with behavioural theory. LiTransMC, fine-tuned using parameter efficient and loss masking strategy, achieved a weighted F1 score of 0.6845 and a Jensen-Shannon Divergence of 0.000245, surpassing both untuned local models and larger proprietary systems, including GPT-4o with advanced persona inference and embedding-based loading, while also outperforming classical mode choice methods such as discrete choice models and machine learning classifiers for the same dataset. This dual improvement, i.e., high instant-level accuracy and near-perfect distributional calibration, demonstrates the feasibility of creating specialist, locally deployable LLMs that integrate prediction and interpretability. Through combining structured behavioural prediction with natural language reasoning, this work unlocks the potential for conversational, multi-task transport models capable of supporting agent-based simulations, policy testing, and behavioural insight generation. These findings establish a pathway for transforming general purpose LLMs into specialized, explainable tools for transportation research and policy formulation, while maintaining privacy, reducing cost, and broadening access through local deployment.
@@
Casual Large Language Models (LLMs), Travel Behavioural Modelling, Open Source LLMs in Transportation, Mode Choice Modelling. Fine-tuning
@@ 1 Introduction
Traditional transportation modelling methods, such as discrete choice models, rely on numerical inputs and structured features to represent traveller preferences. While these methods (e.g., the classic utility-maximization framework of Ben-Akiva and Lerman (Ben-Akiva and Lerman, 1985) have been effective, they often struggle to capture the rich contextual nuances and complex interdependencies of real world travel behaviour. Recent advances in large language models (LLMs) offer a new way to process natural language descriptions of travel scenarios. LLMs can incorporate qualitative factors and domain knowledge that traditional models might miss. Over the past couple of years, researchers have started exploring LLMs for various transportation applications. Early foundational studies showed that general purpose LLMs like GPT-3/43/43 / 4 can leverage their reasoning ability for mobility related predictions. For instance, Wang et al. (2023) reframed human mobility forecasting as a language modelling task, demonstrating that GPT-based models can outperform state-of-the-art time series methods in certain scenarios. Such results hint at a potential paradigm shift from purely data-driven numerical models to LLM-driven approaches that reason about travel behaviour in a human like manner.
@@
The recent surge of interest in generative artificial intelligence has seen LLMs applied across a widening spectrum of domains. These models are trained on vast corpora of text and demonstrate emergent abilities in contextual understanding, multi‑modal data fusion and reasoning, offering a new way for processing complex, human centred information (Zhang et al., 2025). In transportation research, early commentaries by Mahmud et al. (2025) highlighted how LLMs could support intelligent transportation systems (ITS) by predicting traffic flow, detecting vehicles and assisting autonomous driving. Subsequent surveys and frameworks by Nie et al. (2025) categorize the roles of LLMs as information processors, knowledge encoders, component generators and decision facilitators, while Zhang et al. (2025, 2024) illustrate their potential for forecasting mobility time series, integrating unstructured textual inputs and enabling human–machine interaction. At the same time, researchers caution that proprietary LLMs carry high computational costs, and privacy risks; open‑source models and published fine‑tuned derivatives are recommended to ensure transparency, reproducibility and reduced cost (Zhang et al., 2024).
@@
Driven by these opportunities, several studies have begun to explore LLMs for mobility analysis. Wang et al. (2023) introduced LLM‑Mob, a framework that re‑organizes mobility data into historical and context stays and designs context inclusive prompts so that a general purpose GPT‑3.5 model can predict a person’s next location. Their results showed that careful prompt engineering enables LLMs to capture both long‑ and short‑term dependencies and produce accurate, interpretable predictions. Guo et al. (2024) proposed xTP‑LLM for traffic flow prediction, converting multi‑modal traffic data into natural language descriptions and fine‑tuning a LLaMA‑based model with instruction tuning. Their approach achieved competitive accuracy while offering intuitive textual explanations, marking one of the first applications of LLMs to spatio‑temporal traffic forecasting. Other works leverage LLMs as world simulators and assistants Li et al. (2024) presented MobAgent, an agent‑based framework that uses GPT‑4 and DeepSeek APIs to extract mobility patterns and recursively reason about individual motivations in order to generate realistic, personalized travel diaries. These generative applications illustrate the promise of LLMs for producing human‑centred narratives,
@@
More closely related to travel behaviour and mode choice prediction modelling, Mo et al. (2023a) proposed using LLMs to predict travel mode choice in a zero‑shot setting. They design prompts that include task description, individual attributes and domain knowledge and ask ChatGpt 3.5 turbo to predict a commuter’s mode choice without training data. Their results reveal that LLM based predictions can attain accuracy similar to classical discrete choice models and machine‑learning classifiers, and that LLMs can articulate the reasoning behind decisions. However, the authors also observe occasional logical violations and hallucinations in the generated explanations. Liu et al. (2025a) builds on this work and proposes a framework for alignment using a persona‑based embedding learning approach through ChatGPT 4o. They infer personas from socio‑demographic groups through representation learning and condition prompts on these personas. This persona loading significantly improves the match between LLM predictions and human travel choices while remaining computationally tractable. Despite these advances, both studies rely on general‑purpose proprietary models accessed via APIs. No fine‑tuned models specialized for travel choice exist, and there is little discussion of privacy or deployment cost. Moreover, they do not systematically analyze how the textual reasoning aligns with behavioural theory beyond anecdotal observations.
@@
Outside of travel behaviour modelling, some efforts explore the potential of fine-tuned LLMs on transportation safety applications. For example, Zhang et al. (2024) fine‑tuned a LLaMA‑7B model on the TrafficSafety‑2K dataset to produce TrafficSafetyGPT, which outperformed the base LLaMA on transportation safety tasks. The work that comes closest to our own is by Bhandari et al. (2024), who also fine-tuned an open-source model for transport behaviour. In their study, they fine-tuned a Llama-2-7B model on household travel survey data, specifically training it to generate synthetic multi-day travel diaries. Their model learned the sequential patterns of human mobility, producing entire schedules of activities and trips. The crucial distinction lies in the nature of the task and the research objective. Their work addresses a complex, generative challenge of sequence creation, aiming to replicate the statistical distribution of entire travel patterns for applications like large-scale agent-based simulations. Our research, conversely, tackles a targeted predictive challenge, determining the transport mode for a single trip given a specific context. By isolating the mode choice decision, our approach allows for direct performance validation against established discrete choice models and creates a specialized tool for scenario-based policy analysis.
@@
The lack of fine‑tuned LLMs for mode choice prediction indicates a significant research gap. A comprehensive review also urges researchers to prefer open‑source models and to publish fine‑tuned derivatives to facilitate reproducibility (Zhang et al., 2024). Despite this growing interest, research on LLMs in transportation remains limited. Most studies have focused on specific use cases or a single LLM model or framework, restricting the generalizability of findings across diverse contexts. High computational costs often hinder research efforts, while sensitive or proprietary datasets remain inaccessible to many researchers. Data privacy concerns are particularly significant when dealing with personal mobility data, which highlights the need for a comprehensive testing framework. Such framework should evaluate open-access, computationally feasible LLMs, detailing their potential and limitations across various applications.
@@
Building on these insights, our study advances travel choice modelling with three key contributions. First, we provide the most comprehensive evaluation to date of causal, open‑access LLMs for mode choice prediction. We benchmark eleven models (1–12B parameters) across three stated and revealed preference datasets, testing 396 model configurations and generating over 79,000 synthetic predictions. Second, we introduce LiTransMC, the first fine‑tuned causal LLM specialized for mode choice classification. By applying parameter‑efficient low‑rank adaptation (LoRA) with loss masking, we adapt a 12B parameter model using a modest corpus of training examples, achieving superior weighted F1 scores and distributional calibration compared with untuned local models and proprietary GPT‑4o alternatives. Third, beyond predictive accuracy, we develop a systematic framework for analysing LLM reasoning. Using BERTopic and a new Explanation Strength Index, we quantify how LLMs articulate decision factors and align them with behavioural theory.
@@ 2 Methods
The methods are presented in four subsections, detailing the overall experimental design, localized LLM and server setup, evaluation framework, and model finetuning process.
@@ 2.1 Experiment Design
To evaluate the capacity of causal LLMs to model the complex interplay between commuter sociodemographics, trip attributes, and mode choice, and to examine the factors that influence the model’s reasoning and predictive powers, the experiment was structured across multiple layers of variation. These included differences in foundational model selection and size, training and testing datasets, learning strategies, prompting styles, and model configurations, as shown in Figure 1. Eleven foundational models from six different providers were selected, representing the most prominent small- to mid-scale open-access models capable of running efficiently on high-end, user-level local machines. Details of the selected models are provided in Table 1.
@@
As part of the experimental design, three distinct datasets were selected to evaluate the capacity of casual LLMs to predict mode choice and generate corresponding reasoning. The selection aimed to ensure diversity in both survey types, covering stated preference (SP) and revealed preference (RP) approaches, and data accessibility, including both open and closed data sources. This structure was intended to assess how variations in survey methodology and data availability might influence model performance, including the potential effect of prior model exposure to publicly available datasets.
@@
To represent the open-access SP dataset, the widely used Swissmetro survey was selected Bierlaire et al. (2001). For the closed-access SP dataset, a recent experiment conducted in the Brightwater community was utilized. Brightwater is a master-planned, mixed-use development spanning 72 acres of waterfront land in Port Credit, Mississauga, Ontario, Canada. The community is expected to be completed by 2029 and will include over 2,900 residential units and approximately 300,000 square feet of commercial space, with the first residents occupying the space in 2023.
The associated data collection campaign for this study was conducted from October 31 to November 20, 2022. A total of 159 future residents completed the survey, which gathered key sociodemographic information, attitudinal and perception indicators, as well as scenario-based travel responses from Brightwater to major neighbouring destinations across varying seasons and weather conditions.
@@
To represent open-access RP data, the London Passenger Mode Choice (LPMC) dataset was selected (Hillel et al., 2018). This dataset was chosen due to its broad application in mode choice research and its distinctive structure. Unlike most RP surveys, the LPMC dataset includes both the observed choices and the corresponding availability sets of travel alternatives, allowing for the construction of situational and comprehensive choice sets that can be passed to the LLMs for mode choice prediction and reasoning.
@@
For each dataset, a sample of 100 respondents along with their associated scenarios or trips, was selected to construct the training dataset. This training set served as the source pool for generating few-shot examples used in the LLM prompting process. An additional random sample of 200 observations was drawn from the remaining data and used as the testing dataset. This approach ensures that the testing sample size remains consistent across all datasets, while the total size of the training dataset varies depending on the number of scenarios or trips reported per respondent in each survey, with a total of 100 respondents.
@@
To evaluate the most effective learning approach, three different shot types were tested: zero-shot, few-shot with random sampling, and few-shot with targeted, similarity based sampling. In the zero-shot experiments, the commuter’s sociodemographic profile, situational trip attributes, and a contextual description of the travel scenario were provided to the LLM, which acted as a synthetic commuter tasked with selecting the most appropriate mode of transport for the given context. In this setup, the model receives no examples from the original survey, relying solely on its internal knowledge to infer the relationship between individual characteristics, trip attributes, and mode choice. Thus, it makes decisions based entirely on the presented context and its generalised understanding, without exposure to any observed examples of human decision making from the training data. In contrast, the few-shot approach involves providing selected examples from the training dataset to the LLM in order to support its understanding of the nuanced relationships between sociodemographic profiles, trip attributes, and mode choice decisions. Within this few-shot learning modality, two sampling strategies were employed to select examples from the training pool.
@@
For the Swissmetro dataset, the weights are defined as:
@@
Before finalizing the current similarity framework, several alternative approaches were explored, including cosine similarity and cross-encoder embeddings for full-text comparison. While cosine similarity is widely used in natural language processing for comparing semantic vectors, it underperforms when applied to numeric survey responses, such as travel time or cost. This is because cosine similarity measures the angle between vectors rather than their magnitude, making it insensitive to absolute differences that are critical for behavioural decisions in mode choice modelling.
For instance, consider a test instance with a scaled travel cost of 0.81 (representing 240 CHF). Two candidate training examples have scaled values of 0.8 (237 CHF) and 0.20 (65 CHF), respectively. Under cosine similarity, both vectors yield high similarity scores (e.g., >0.98>0.98> 0.98) because their directions are nearly aligned, despite the second case representing a cost difference of 175 CHF. In contrast, using inverse Euclidean distance, the similarity scores are more behaviourally meaningful:
@@
For categorical and ordinal variables, cosine similarity similarly fails to capture behavioural meaning, as it treats differences between categories like “leisure” and “business”, or between income levels 2 and 3, as abstract angular dissimilarities rather than ordered or discrete groupings. In contrast, the adopted framework uses exact matching for nominal variables and ordinal proximity scoring for ordered variables as described previously in Equation 3, thereby better reflecting real-world perception thresholds (e.g., adjacent age or income groups are more similar than distant ones).
[…]

### v2

@@ Abstract
This study investigates the adoption of open-access, locally deployable causal large language models (LLMs) for travel mode choice prediction and introduces LiTransMC, the first fine-tuned causal LLM developed for this task. We systematically benchmark eleven open-access LLMs (1-12B parameters) across three stated and revealed preference datasets, testing 396 configurations and generating over 79,000 mode choice decisions. Beyond predictive accuracy, we evaluate models generated reasoning using BERTopic for topic modelling and a novel Explanation Strength Index, providing the first structured analysis of how LLMs articulate decision factors in alignment with behavioural theory. LiTransMC, fine-tuned using parameter efficient and loss masking strategy, achieved a weighted F1 score of 0.6845 and a Jensen-Shannon Divergence of 0.000245, surpassing both untuned local models and larger proprietary systems, including GPT-4o with advanced persona inference and embedding-based loading, while also outperforming classical mode choice methods such as discrete choice models and machine learning classifiers for the same dataset. This dual improvement, i.e., high instant-level accuracy and near-perfect distributional calibration, demonstrates the feasibility of creating specialist, locally deployable LLMs that integrate prediction and interpretability. Through combining structured behavioural prediction with natural language reasoning, this work unlocks the potential for conversational, multi-task transport models capable of supporting agent-based simulations, policy testing, and behavioural insight generation. These findings establish a pathway for transforming general purpose LLMs into specialized and explainable tools for transportation research and policy formulation, while maintaining privacy, reducing cost, and broadening access through local deployment.
@@
Casual Large Language Models (LLMs), Travel Behavioural Modelling, Open Source LLMs in Transportation, Mode Choice Modelling. LLMs Fine-tuning
@@ 1 Introduction
Traditional transportation modelling has evolved through several methodological advances, each building upon the previous generation, yet leaving some important gaps. The random utility maximization framework introduced by McFadden (McFadden, 1974) and formalized in Ben-Akiva and Lerman’s seminal textbook (Ben-Akiva and Lerman, 1985) established discrete choice models as the foundation of travel demand analysis. These models provided behavioural interpretability, but relied on restrictive assumptions such as independence of irrelevant alternatives (IIA) and linear-in-parameters utility functions, which constrained their ability to capture complex substitution patterns. To overcome some of these restrictions, simulation-based estimation methods were introduced (Train, 2009), and hybrid choice models incorporated attitudinal and perceptual constructs (Ben-Akiva et al., 2002). While these extensions enhanced behavioural realism, they still depended on structured survey variables and struggled to integrate unstructured or contextual information that influences traveller decisions.
Parallel explorations in machine learning (ML) offered another direction. Neural networks and related classifiers (Shmueli et al., 1996; Hensher and Ton, 2000) demonstrated higher predictive accuracy by capturing non-linearities and interactions in addition to their ability to incorporate other ubiquitous data, like gps trajectories and WiFi signals for choice modelling. However, ML models typically sacrificed interpretability and departed from established behavioural theory. In practice, this meant that while econometric models could explain why a decision was made but sometimes lacked predictive power, machine learning models could predict choices without offering transparent behavioural insights. Moreover, the practical execution of travel surveys is increasingly hindered by declining response rates and recruitment difficulties, which undermine the representativeness of samples and introduce nonresponse bias in ways that are difficult to correct (Wittwer et al., 2024; Wang et al., 2023b; Svaboe et al., 2024).
This growing debate between behavioural realism and predictive power combined with lower participation rates sets the stage for advances from Large Language Models (LLMs), which provide the capability to process natural language, incorporate qualitative context, and articulate reasoning in ways that earlier models could not. Additionally, their capacity for few-shot generalization enables the inference of behavioural regularities from limited examples, which supports the creation of data-efficient digital representations of respondents that capture contextual and cognitive diversity beyond what conventional surveys can offer. In natural language processing, early recurrent and LSTM-based models (Hochreiter and Schmidhuber, 1997; Mikolov et al., 2010) improved over statistical n-grams, but the real breakthrough came with the transformer architecture (Vaswani et al., 2017). This enabled large-scale pretraining, exemplified by BERT’s bidirectional encoder (Devlin et al., 2019) and the GPT series of autoregressive decoders (Brown et al., 2020). Subsequent work on scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), retrieval-augmented generation (Borgeaud et al., 2022), and alignment methods such as the Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) shaped the trajectory toward today’s frontier models, including closed systems like the proprietary GPT series (Achiam et al., 2023) and open-source alternatives such as LLaMA (Touvron et al., 2023).
@@
The recent surge of interest in generative artificial intelligence has seen LLMs applied across a widening spectrum of domains. These models are trained on vast corpora of text and demonstrate emergent abilities in contextual understanding, multi‑modal data fusion and reasoning, offering a new way for processing complex, human centred information (Zhang et al., 2025). In transportation research, early commentaries by Mahmud et al. (2025) highlighted how LLMs could support intelligent transportation systems (ITS) by predicting traffic flow, detecting vehicles and assisting autonomous driving. Subsequent surveys and frameworks by Nie et al. (2025) categorize the roles of LLMs as information processors, knowledge encoders, component generators and decision facilitators, while Zhang et al. (2025, 2024) illustrate their potential for forecasting mobility time series, where they integrate unstructured textual inputs and enable human–machine interaction. At the same time, researchers caution that proprietary LLMs carry high computational costs, and privacy risks; open‑source models and published fine‑tuned derivatives are recommended to ensure transparency, reproducibility and reduced cost (Zhang et al., 2024).
@@
Driven by these opportunities, several studies have begun to explore LLMs for mobility analysis. Wang et al. (2023a) introduced LLM‑Mob, a framework that reorganizes mobility data into historical and context stays and designs context inclusive prompts so that a general purpose GPT‑3.5 model can predict a person’s next location. Their results showed that careful prompt engineering enables LLMs to capture both long‑ and short‑term dependencies and produce accurate, interpretable predictions. Guo et al. (2024) proposed xTP‑LLM for traffic flow prediction, converting multi‑modal traffic data into natural language descriptions and fine‑tuning a LLaMA‑based model with instruction tuning. Their approach achieved competitive accuracy while offering intuitive textual explanations, marking one of the first applications of LLMs to spatio‑temporal traffic forecasting. Other works leverage LLMs as world simulators and assistants Li et al. (2024) presented MobAgent, an agent‑based framework that uses GPT‑4 and DeepSeek APIs to extract mobility patterns and recursively reason about individual motivations in order to generate realistic, personalised travel diaries. These generative applications highlight the promise of LLMs for producing human‑centred narratives.
@@
More closely related to travel behaviour and mode choice prediction modelling, Mo et al. (2023) proposed using LLMs to predict travel mode choice in a zero‑shot setting. They design prompts that include task description, individual attributes and domain knowledge and ask ChatGPT 3.5 turbo to predict a commuter’s mode choice without any training data. Their results reveal that LLM based predictions can attain accuracy similar to classical discrete choice models and machine‑learning classifiers, and that LLMs can articulate the reasoning behind decisions. However, the authors also observe occasional logical violations and hallucinations in the generated explanations. This highlights how granting LLMs access to open-access datasets and domain-specific research corpora could substantially enhance their capacity to acquire specialised transport knowledge, moving beyond the limitations of general-purpose pre-training. In Liu et al. (2025) work, we do not see a replication of the accuracy at the zero-shot level, but the authors build on this work and propose a framework for alignment using a persona‑based embedding learning approach through ChatGPT 4o for few-shot prompting. They infer personas from socio‑demographic groups through representation learning and condition prompts on these personas. This persona loading significantly improves the match between LLM predictions and human travel choices while remaining computationally tractable. Despite these advances, both studies rely on general‑purpose proprietary models accessed via APIs. No fine‑tuned models specialized for travel choice exist, and there is little discussion of privacy or deployment cost. Moreover, these studies do not provide a systematic analysis of how the textual reasoning aligns with behavioural theory beyond anecdotal observations.
@@
Outside of travel behaviour modelling, some efforts explore the potential of fine-tuned LLMs on transportation safety applications. For example, Zhang et al. (2024) fine‑tuned the LlaMA‑7B model on the TrafficSafety‑2K dataset to produce TrafficSafetyGPT, which outperformed the base LlaMA on transportation safety tasks. The work that comes closest to our own is by Bhandari et al. (2024), who also fine-tuned an open-source model for transport behaviour. In their study, they fine-tuned a Llama-2-7B model on household travel survey data, specifically training it to generate synthetic multi-day travel diaries. The model learned the sequential patterns of human mobility, producing entire schedules of activities and trips. The crucial distinction lies in the nature of the task and the research objective. This work addresses a complex, generative challenge of sequence creation, aiming to replicate the statistical distribution of entire travel patterns for applications like large-scale agent-based simulations. Our research, conversely, tackles a targeted predictive challenge, determining the travel mode for a single trip given a specific context. By isolating the mode choice decision, our approach allows for direct performance validation against established discrete choice models and creates a specialized tool for scenario-based policy analysis.
@@
The lack of fine‑tuned LLMs for travel mode choice prediction indicates a significant research gap. A comprehensive review also urges researchers to prefer open‑source models and to publish fine‑tuned derivatives to facilitate reproducibility (Zhang et al., 2024). Despite a growing interest, research on LLMs in transportation remains limited. Most studies have focused on specific use cases or a single LLM model or framework, restricting the generalizability of findings across diverse contexts. High computational costs often hinder research efforts, while sensitive or proprietary datasets remain inaccessible to many researchers. Data privacy concerns are particularly significant when dealing with personal mobility data, which highlights the need for a comprehensive testing framework. Such a framework should evaluate open-access, computationally feasible LLMs, detailing their potential and limitations across various applications.
Building on these insights, our study advances travel choice modelling with three key contributions.
  • First, we provide the most comprehensive evaluation to date of causal, open‑access LLMs for mode choice prediction. We benchmark eleven open-access models (1–12B parameters) across three stated and revealed preference datasets (two open-access and one closed-access), testing 396 model configurations and generating over 79,000 mode choice predictions.
@@
  • Second, we introduce LiTransMC, to the best of our knowledge, the first fine‑tuned causal LLM specialized for mode choice classification. By applying parameter‑efficient low‑rank adaptation (LoRA) with loss masking, we adapt a 12B parameter model using a modest corpus of training examples, achieving superior weighted F1 scores and distributional calibration compared with untuned local models and proprietary GPT‑4o alternatives.
  • Third, beyond predictive accuracy, we develop a systematic framework for analyzing LLM reasoning. Using BERTopic and a new Explanation Strength Index, we quantify how LLMs articulate decision factors and align them with behavioural theory.
The rest of the paper is organized as follows. Section 2 details the methodological framework, including the experimental design, dataset selection, learning strategies, and local LLM setup. Section 3 presents and discusses the results, focusing on predictive accuracy, reasoning performance, and the effects of fine-tuning. Section 4 outlines key recommendations and implications for transport behaviour research and practice, and Section 5 concludes the paper with final reflections and directions for future work.
@@ 2 Methods
Here we present the overall experimental design, localized LLM and server setup, evaluation framework, and LLM model finetuning process.
@@ 2.1 Experiment Design
To evaluate the ability of causal LLMs to model the complex relationship between commuters sociodemographics, trip attributes, and mode choice, and to examine the factors that influence the model’s predictive and reasoning powers, the experiment was developed with a wide range of variations. These included varying the chosen foundational models, the size of the model parameters, training and testing datasets, learning strategies, prompting styles, and the model’s temperature configuration. Figure 1 outlines the overall experiment configuration and the variation in the experiment variables.
@@
2.1.1 Foundational Models
Eleven foundational models from six different providers were selected, representing the most prominent small- to mid-scale open-access models. The models were chosen so that they can be efficiently deployed on high-end, consumer-level local desktop machines. Details of the selected models are provided in Table 1.
@@ 2.1.2 Datasets
Three distinct datasets were selected to evaluate the ability of causal LLMs to predict commuters mode choice and generate corresponding reasoning. The selection aimed to ensure diversity in both survey types and data accessibility, covering stated preference (SP) and revealed preference (RP) surveys, with open and closed data sources. This structure was intended to assess how variations in survey methodology and data availability might influence model performance, including the potential effect of prior model exposure to publicly available datasets.
@@
To represent the open-access SP dataset, the widely used Swissmetro survey was selected (Bierlaire et al., 2001). For the closed-access SP dataset, a recent experiment conducted in the Brightwater community was utilized. Brightwater is a master-planned, mixed-use development spanning 72 acres of waterfront land in Port Credit, Mississauga, Ontario, Canada. The community is expected to be completed by 2029 and will include over 2,900 residential units and approximately 300,000 square feet of commercial space, with the first residents occupying the space in 2023. The associated data collection campaign for this study was conducted from October 31 to November 20, 2022. A total of 159 future residents completed the survey, which gathered key sociodemographic information, attitudinal and perception indicators, as well as scenario-based travel responses from Brightwater to major neighbouring destinations across varying seasons and weather conditions.
@@
To represent open-access RP data, the London Passenger Mode Choice (LPMC) dataset was selected (Hillel et al., 2018). This dataset was chosen due to its broad application in mode choice research and its distinctive structure. Unlike most RP surveys, the LPMC dataset includes both the observed choices and the corresponding availability sets of travel alternatives, allowing to construct a situational choice sets.
@@
For each dataset, a sample of all scenarios or trips for a 100 respondents was selected to construct the training dataset. This training set served as the source pool for generating few-shot examples used in the LLM prompting process. An additional random sample of 200 observations was drawn from the remaining data and used as the testing dataset. This approach ensures that the testing sample size remains consistent across all datasets, while the total size of the training dataset varies depending on the number of scenarios or trips reported per respondent in each survey for a 100 respondent profiles.
2.1.3 Learning Approach
@@
To evaluate the most effective learning approach, three different shot types were tested: zero-shot, few-shot with random sampling, and few-shot with targeted, similarity based sampling. In the zero-shot experiments, the commuter’s sociodemographic profile, situational trip attributes, and a contextual description of the travel scenario were provided to the LLM, which acted as a synthetic commuter tasked with selecting the most appropriate mode of transport for the given context. In this setup, the model receives no examples from the original survey, relying solely on its internal knowledge to infer the relationship between individual characteristics, trip attributes, and mode choice. Thus, it makes decisions based entirely on the presented context and its generalized understanding, without exposure to any observed examples of human decision making from the training data. On the other hand, the few-shot approach involves providing selected examples from the training dataset to the LLM in order to support its understanding of the relationships between sociodemographic profiles, trip attributes, and mode choice decisions. Within this few-shot learning modality, two sampling strategies were employed to select examples from the training pool.
@@
For example, in the Swissmetro dataset, the weights are defined as:
@@
Before finalizing the current similarity framework, several alternative approaches were explored, including cosine similarity and cross-encoder embeddings for full-text comparison. While cosine similarity is widely used in natural language processing for comparing semantic vectors, and in previous studies to select the few-shot examples in mode choice experiments (Liu et al., 2025), it underperforms when applied to numeric survey responses, such as travel time or cost. This is because cosine similarity measures the angle between vectors rather than their magnitude, making it insensitive to absolute differences that are critical for behavioural decisions in mode choice modelling.
For instance, if we consider a test instance with a scaled travel cost of 0.81 (representing 240 CHF). Two candidate training examples have scaled values of 0.8 (237 CHF) and 0.20 (65 CHF), respectively. Under cosine similarity, both vectors yield high similarity scores (e.g., >0.98>0.98) because their directions are aligned, despite the second case representing a cost difference of 175 CHF. In contrast, using inverse Euclidean distance, the similarity scores are more behaviourally meaningful:
@@
For categorical and ordinal variables, cosine similarity similarly fails to capture behavioural meaning, as it treats differences between categories like “leisure” and “business”, or between income levels 2 and 3, as abstract angular dissimilarities rather than ordered or discrete groupings. The adopted framework, on the other hand, uses exact matching for nominal variables and ordinal proximity scoring for ordered variables as described previously in Equation 2, thereby better reflecting real-world perception thresholds (e.g., adjacent age or income groups are more similar than distant ones).
Another element to consider lies in the design of the prompt construction pipeline. Under zero-shot and random few-shot configurations, survey responses are pre-processed by converting structured attributes (both categorical and continuous) into full textual descriptions. These narrative prompts are stored in a textual database and retrieved directly for LLM calls, minimizing computational cost during runtime. However, for the targeted similarity-based setup, the similarity computation is performed on the original structured data (prior to textual transformation). For each test instance, the top-$k$ most similar training samples are first identified, and their corresponding prompts are dynamically constructed at runtime. This ensures that the similarity is calculated without semantic noise from full-text embeddings.
2.1.4 Prompting Style